# Stabilizing Non-Abelian Topological Order against Heralded Noise via Local Lindbladian Dynamics

### v1

@@ IV.2 Error Correction Protocol
Recall that a $Z_{k_{c}}$ ($c=r,g,b)$ error on a single edge anticommutes with the two $A_{p_{c}}$ plaquette operators sharing that edge; thus, strings of $Z$ errors create pairs of Abelian anyons that always fuse back to the vacuum $A_{p_{c}}=+1$ when brought together. In the previous section, we observed that new $Z$ flags and plaquette defects emerge as a result of vertex correction processes. The flags that herald these new defects are only raised once the $X$ flags in the vicinity of the plaquette are removed. If we instead measure the plaquette before the complete set of $Z$ flags are available, the defects may get pushed in a direction incompatible with that of their paired partner. We avoid this by requiring that the $X$ flags on qubits enclosed inside a plaquette $p_{c}$ must be removed before measuring $A_{p_{c}}$. This ensures that the potential defect $A_{p_{c}}=-1$ will be perfectly heralded with respect to its partner defect.
@@ IV.3 Simulation using Non-Pauli stabilizer tableau
=$\begin{split}\mathcal{L}&{}_{C_{c}}=\gamma_{x}\sum_{\begin{subarray}{c}v_{c},s=\pm 1\\ j\in^{*}v_{c},\mathbf{m^{z}}\in\{0,1\}^{\otimes 4}\end{subarray}}\mathcal{L}^{{D_{4}}^{\text{X-leaf}}}_{v_{c},j,s,\mathbf{m^{z}}}\ +\ \gamma_{z}\sum_{\begin{subarray}{c}p_{c},s=\pm 1\\ j\in\partial p_{c}\end{subarray}}\mathcal{L}^{{D_{4}}^{\text{Z-leaf}}}_{p_{c},j,s}\\ &+\gamma_{x}\sum_{\begin{subarray}{c}v_{c}\in\text{sublattice-1},\\ s=\pm 1,\mathbf{m^{x}}\in\{0,1\}^{\otimes 4}\\ ,\mathbf{m^{z}}\in\{0,1\}^{\otimes 6}\end{subarray}}\mathcal{L}^{{D_{4}}^{\text{X-loop}}}_{v_{c},s,\mathbf{m^{x}},\mathbf{m^{z}}}+\gamma_{z}\sum_{\begin{subarray}{c}p_{c},s=\pm 1,\\ \varepsilon\in\mathcal{E}(p_{c}),\mathbf{m^{z}}\end{subarray}}\mathcal{L}^{{D_{4}}^{\text{Z-loop}}}_{p_{c},s,\varepsilon,\mathbf{m^{z}}}\ .\end{split}$ (34)
@@
In our protocol, the flags undergo classical dynamics, and can thus be simulated using standard Monte Carlo techniques (see, [103]). However, as mentioned before, one complication of the above correction protocol, relative to that required for the Toric code, is that the dynamics is no longer diagonal in the quasi-stabilizer basis. This is because applying Pauli-$X$ operators creates a superposition of different configurations of Abelian anyons. Hence, simulating the full dynamics of the quantum system requires additional techniques. Here, we adapt the stabilizer tableau method [92, 93] and describe how an appropriate generalization allows us to simulate the dynamics of the quasi-stabilizer defects. Specifically, this method allows us to simulate individual measurement trajectories of the quasi-stabilizers, which we then average over independently generated measurement trajectories using a Monte Carlo approach to compute physical observables (see Appendix. A). The key difference in our approach, relative to usual stabilizer tableaux, is that the quasi-stabilizers defining the $D_{4}$ state are non-commuting operators that are not elements of the Pauli group. We outline how this complication can be treated here, with further details provided in Appendix C. As before, we slightly abuse the notation and refer to $A_{p}$ and $B_{v}$ as stabilizers whenever it is clear from the context.
@@
Figure 10: The modification of plaquette stabilizer $A_{p}$ after repeated application of $X$ on its interior edges. ($\rightarrow$ b): applying $X$ to a green edge $e_{1}$ appends the product $Z_{e_{0}}Z_{e_{2}}$ of Pauli $Z$ operators on the two adjacent red edges to the original plaquette operator. ($\rightarrow$ c): applying $X$ to a red edge $e_{2}$ appends $Z_{e_{1}}Z_{e_{3}}$ to the plaquette operator along with a minus sign, due to anticommutation with the existing $Z_{e_{2}}$ from the previous step. Updating vertex stabilizers, or updating plaquette stabilizers after applying $Z$ to an external edge, changes the resulting stabilizers only up to a sign.
@@
Next, to determine the effect of measurements we use the fact that at any point during the simulation, the instantaneous stabilizers have the form {±$\{\pm B_{v},\pm\prod_{k,p}Z_{k}A_{p}\}$. These operators either commute or anticommute with the operators $B_{v}$ and $A_{p}$ (when acting on the state at that instant of time) that we measure, so standard stabilizer tableau techniques can be used to simulate these measurement processes. $B_{v}$ commutes with all instantaneous stabilizers, and always has a definite measurement outcome given by the sign of the associated instantaneous vertex stabilizer. $A_{p}$, on the other hand, may either commute or anti-commute with a given instantaneous stabilizer. When it fails to commute, the instantaneous stabilizers must be modified in a manner consistent with the probabilistic measurement outcome (see Appendix C). If $A_{p}$ commutes with all instantaneous stabilizers, the measurement outcome is deterministic and the set of stabilizers need not be modified. However, it can be the case that $A_{p}$ commutes with all instantaneous stabilizers, but is not among the instantaneous list of stabilizer generators. In this case, this deterministic measurement outcome of $A_{p}$ must be determined from the existing stabilizer generators; we describe how to do this in in Appendix C and also provide additional details about encoding the stabilizers in tableaux that can be efficiently simulated.
@@
Figure 11: Time evolution for the $D_{4}$ model on three $L\times L$ sized interpenetrating honeycomb lattices at three representative points in the $\gamma_{x}-\gamma_{z}$ phase diagram (see Fig. 1(b)). Plots show the densities of $X$ and $Z$ flags, stabilizer defect densities $n_{d}^{B}$ and $n_{d}^{A}$, and the logical error rate $1-F_{|\widetilde{0}\rangle_{L}(|\widetilde{+}\rangle_{L})}$ (see Eq. (41)) as a function of time. Time is measured in units of inverse erasure noise rate, set to $\eta=1$. Note that the rate of correction processes relative to the noise is much higher compared to the Toric code (see Fig. 5), leading to a comparatively shorter extent of the X axis for similar number of Monte Carlo sweeps. The flags are initialized to $n^{f}_{X}=n^{f}_{Z}=0$ and the initial state of qubits is set to $|\widetilde{0}\rangle_{L}$, except when calculating $1-F_{|\widetilde{+}\rangle_{L}}$, where it is initialized in the |+~$|\widetilde{+}\rangle_{L}$ state. Each data point is obtained by averaging over $500$ independent Monte Carlo trajectories. (a) Absorbing phase ($\gamma_{x}=4,\gamma_{z}=500$): when the $X$ correction rate is small, $X$ flags proliferate, causing the density of vertex defects to saturate to $n^{d}_{B}=1/2$. Because the $Z$ correction operations in Eqs. (31), (32) can only be carried out when $X$ flags are absent on all edges inside a plaquette, they are necessarily ineffective when $n^{f}_{X}=1$. Thus, $Z$ flags also enter the absorbing state, irrespective of the $Z$ correction rate, causing plaquette defects to saturate to $n^{d}_{A}=1/2$. In this region all logical information is lost, as witnessed by $F\rightarrow 0$. (b) Partially $X$-active phase ($\gamma_{x}=14,\gamma_{z}=500$): densities of $X$ flags and vertex defects saturate to small values, since the $X$ correction rate exceeds the corresponding Toric code threshold. However, even the modest steady-state $X$ flag density suffices to make the effective $Z$ correction rate unable to prevent the $Z$ flags and plaquette defects from saturating to their absorbing state values. This phase can only protect classical information, as seen by perfect recovery of $|\widetilde{0}\rangle_{L}$ state (magenta curve) but failure to preserve |+~$|\widetilde{+}\rangle_{L}$ state (cyan curve). (c) Active state ($\gamma_{x}=35,\gamma_{z}=500$): increasing the $X$ correction rate further decreases the long-time value of $n^{f}_{X}$ (and of $n^{d}_{B})$, allowing $Z$ correction to become sufficiently effective to stabilize $n^{f}_{Z}$ and $n^{d}_{A}$ densities at small values. In this phase, the logical error probability is close to 00 witnessed by $(1-F)\approx 0$ for both $|\widetilde{0}\rangle_{L}$ and |+~$|\widetilde{+}\rangle_{L}$ initial states, indicating that logical information is preserved out to the longest times accessible numerically. Note that our simulation protocol in the active phase is numerically less intensive compared to the absorbing phases since most $A$ stabilizers are unmodified due to low density of defects. Hence, the system sizes shown in (a) and (b) ($L=24$) are smaller than those $L=30$ accessible in the active phase (c).
@@ V.1 Dynamical phase diagram of flags
Fig. 11 shows the combined evolution of flag densities, stabilizer defect densities, and logical fidelity at a representative parameter value within each distinct steady-state phase shown in Fig. 1(b). We see that in each case, the dynamics of stabilizer defects closely tracks that of the flags, with both steady-state values and the time-scales at which these reach their long-time values closely correlated. Here, the densities of vertex and plaquette stabilizer defects are defined respectively as
@@ V.2 Diagnostics for topologically ordered steady-states
Let us begin with the error-correction metric. The simultaneous $+1$ eigenspace of the $A_{p}$ and $B_{v}$ operators of the $D_{4}$ model is $22$-fold degenerate on the torus, reflecting the number of distinct superselection sectors (anyons) of $D_{4}$ TO. The states in this subspace can be distinguished based on the values of certain non-contractible loop operators. One choice is to label the $22$ ground states according to their eigenvalues under the non-contractible Abelian loop operators, given by $\mathcal{Z}^{c}_{H/V}=\prod_{k\in l_{H/V}}Z_{k}=\pm 1$ where $l_{H}(l_{V})$ are horizontal (vertical) loops on the dual of the $c$ colored honeycomb lattice, as shown in Fig 12. We term these the “logical-$Z$” operators. While there are $2^{6}$ different choices for the eigenvalues of these logical operators, only $22$ of them are consistent with $A_{p}=B_{v}=1$ for all $p,v$ [64]. The operators that switch between the allowed logical-$Z$ eigenstates are the non-contractible string operators associated with transporting non-Abelian anyons around the torus. We refer to these as the “logical-$X$” operators, $\mathcal{X}^{c}_{H},\mathcal{X}^{c}_{V}$. As shown in Fig 12, these non-Abelian string operators consist of a string of $X$ operators decorated with non-local $CZ$ gates. A second basis choice for the $22$ locally indistinguishable ground states is by their values under the vertical $\mathcal{Z}_{V}$ and $\mathcal{X}_{V}$ logical operators.
In order to analyze the fate of the encoded logical information in the ground-spae of the $D_{4}$ model under the noise and correction dynamics, it will be useful to consider two different initial logical states, one in each of these basis choices:
@@
Figure 13: Two way connectivity in the $D_{4}$ active phase: we estimate the time required for noiseless dynamics $\eta=0$ to take the (quasi)-steady-state $\rho$ in the active phase ($\gamma_{x}=35,\gamma_{z}=500$ simulated upto $500$ MC sweeps) back to the pure $D_{4}$ ground state $|\psi_{0}\rangle$. To numerically access the larger system sizes, we estimate the fidelity of the time evolved state based on the flag densities, namely $F_{\text{flag}}(t):=\langle n^{f}_{X,Z}=0|\rho(t)|n^{f}_{X,Z}=0\rangle$, where $|n^{f}_{X,Z}=0\rangle$ is the state where $X$ and $Z$ flags on all sites are unoccupied (see Eq. (42)). (a) $1-F_{\text{flags}}(t)$ evaluated for various system sizes with linear dimension $L$. The data is obtained by averaging over $1000$ independent Monte Carlo realizations. (b) $1-F_{\text{flags}}(t)$ (blue) and $1-F(t)$ of qubit system (orange) are plotted as a function of time for $L=30$. This demonstrates that $F_{\text{flags}}(t)<F_{\text{full}}(t)$ and hence recovery time-scales obtained using flag-fidelities upperbound the recovery time computed in terms of the full fidelity of stabilizer system. (c) Time $\tau$ to achieve $1-F_{\text{flag}}(t)>\varepsilon$ is plotted as a function of system size $L$ for various cutoff values $\varepsilon$. Here, the X-axis is shown on a log-scale. The dotted lines show best fits to the functional form $\alpha\log{(L/L_{0})}$. The fit quality diminishes with increasing $\varepsilon$ since when most trajectories have recovered (i.e $1-F\sim 0$), the fluctuations due to a limited number of samples $N_{r}$ become more pronounced. We note that presence of even single flag in a given trajectory leads to zero fidelity with $|n^{f}_{X,Z}\rangle$ state. While the time-scale for the state to get close to the initial flag-free state in terms of fidelity increases with increasing $L$, the local densities of flags approach zero in time independent of the system size.
Our second diagnostic utilizes the two-way channel connectivity equivalence relation placed on the space of mixed-states [53]: two states $\rho_{1}$ and $\rho_{2}$ belong to the same mixed-state phase if there exist a pair of quasi-local, finite-depth quantum channels $\Sigma_{12},\Sigma_{21}$ such that $\Sigma_{21}(\rho_{1})=\rho_{2}$ and $\Sigma_{12}(\rho_{2})=\rho_{1}$. In our case, the challenge is to determine whether there exists a finite-depth quantum channel that takes the steady-state density matrix $\rho_{ss}$ back to the initial error-free state $\rho_{1}$. We assess this by measuring the time required for the steady-state to evolve back to the error-free state under the Lindbladian Eq. (33) in the absence of noise i.e., under the action of only the error correction terms. Specifically, we examine how the time $t$ at which $\rho_{ss}\sim e^{\mathcal{L}_{\eta=0}t}\rho_{1}$ scales with system size, with $\mathcal{L}_{\eta=0}$ given by Eq. (33) with $\eta=0$. The density matrices $\rho_{1}$ and $\rho_{ss}$ are considered as being connected via a finite-depth quantum channel as long as this time grows at most (poly)-logarithmically with $L$ [53]. Here two mixed states are considered to be close in terms of the fidelity, i.e. $\rho\sim\rho^{\prime}$, if $F(\rho,\rho^{\prime})>1-\varepsilon$ for some $\varepsilon>0$.
We emphasize that this is, in principle, a different (and stronger) criterion than requiring that the logical information encoded in the initial state can be recovered, since our MWPM decoder is highly non-local and need not guarantee the existence of a pair of quasi-local finite-depth two-way channels. Interestingly however, at least in the context of the Toric code, Ref. [53] showed that the two criteria are equivalent in practice.

### v2

@@ IV.2 Error Correction Protocol
Recall that a $Z_{k_{c}}$ ($c=r,g,b)$ error on a single edge anticommutes with the two $A_{p_{c}}$ plaquette operators sharing that edge; thus, strings of $Z$ errors create pairs of Abelian anyons that always fuse back to the vacuum $A_{p_{c}}=+1$ when brought together. In the previous section, we observed that new $Z$ flags and plaquette defects emerge as a result of vertex correction processes. The flags that herald these new defects are only raised once the $X$ flags in the vicinity of the plaquette are removed. If we instead measure the plaquette before the complete set of $Z$ flags is available, the defects may get pushed in a direction incompatible with that of their paired partner. We avoid this by requiring that the $X$ flags on qubits enclosed inside a plaquette $p_{c}$ must be removed before measuring $A_{p_{c}}$. This ensures that the potential defect $A_{p_{c}}=-1$ will be perfectly heralded with respect to its partner defect.
@@ IV.3 Simulation using Non-Pauli stabilizer tableau
$\mathcal{L}_{C_{c}}=\gamma_{x}\sum_{\begin{subarray}{c}v_{c},s=\pm 1\\ j\in^{*}v_{c},\\ \mathbf{m^{z}}\in\{0,1\}^{\otimes 4}\end{subarray}}\mathcal{L}^{{D_{4}}^{\text{X-leaf}}}_{v_{c},j,s,\mathbf{m^{z}}}\ +\ \gamma_{z}\sum_{\begin{subarray}{c}p_{c},s=\pm 1\\ j\in\partial p_{c}\end{subarray}}\mathcal{L}^{{D_{4}}^{\text{Z-leaf}}}_{p_{c},j,s}$ (35)
$+\gamma_{x}\sum_{\begin{subarray}{c}v_{c}\in\text{sublattice-1},\\ s=\pm 1,\\ \mathbf{m^{x}}\in\{0,1\}^{\otimes 4}\\ ,\mathbf{m^{z}}\in\{0,1\}^{\otimes 6}\end{subarray}}\mathcal{L}^{{D_{4}}^{\text{X-loop}}}_{v_{c},s,\mathbf{m^{x}},\mathbf{m^{z}}}+\gamma_{z}\sum_{\begin{subarray}{c}p_{c},\\ s=\pm 1,\\ \varepsilon\in\mathcal{E}(p_{c}),\\ \mathbf{m^{z}}\end{subarray}}\mathcal{L}^{{D_{4}}^{\text{Z-loop}}}_{p_{c},s,\varepsilon,\mathbf{m^{z}}}.$
@@
In our protocol, the flags undergo classical dynamics and can thus be simulated using standard Monte Carlo techniques (see [81]). However, as mentioned before, one complication of the above correction protocol, relative to that required for the Toric code, is that the dynamics is no longer diagonal in the quasi-stabilizer basis. This is because applying Pauli-$X$ operators creates a superposition of different configurations of Abelian anyons. Hence, simulating the full dynamics of the quantum system requires additional techniques. Here, we adapt the stabilizer tableau method [96, 97] and describe how an appropriate generalization allows us to simulate the dynamics of the quasi-stabilizer defects. Specifically, this method allows us to simulate individual measurement trajectories of the quasi-stabilizers, which we then average over independently generated measurement trajectories using a Monte Carlo approach to compute physical observables (see Appendix. A). The key difference in our approach, relative to usual stabilizer tableaux, is that the quasi-stabilizers defining the $D_{4}$ state are non-commuting operators that are not elements of the Pauli group. We outline how this complication can be treated here, with further details provided in Appendix C. As before, we slightly abuse the notation and refer to $A_{p}$ and $B_{v}$ as stabilizers whenever it is clear from the context.
@@
[…]
@@
Next, to determine the effect of measurements, we use the fact that at any point during the simulation, the instantaneous stabilizers have the form $\{\pm B_{v},\pm\prod_{k,p}Z_{k}A_{p}\}$. These operators either commute or anticommute with the operators $B_{v}$ and $A_{p}$ (when acting on the state at that instant of time) that we measure, so standard stabilizer tableau techniques can be used to simulate these measurement processes. $B_{v}$ commutes with all instantaneous stabilizers, and always has a definite measurement outcome given by the sign of the associated instantaneous vertex stabilizer. $A_{p}$, on the other hand, may either commute or anti-commute with a given instantaneous stabilizer. When it fails to commute, the instantaneous stabilizers must be modified in a manner consistent with the probabilistic measurement outcome (see Appendix C). If $A_{p}$ commutes with all instantaneous stabilizers, the measurement outcome is deterministic, and the set of stabilizers need not be modified. However, it can be the case that $A_{p}$ commutes with all instantaneous stabilizers, but is not among the instantaneous list of stabilizer generators. In this case, this deterministic measurement outcome of $A_{p}$ must be determined from the existing stabilizer generators; we describe how to do this in Appendix C and also provide additional details about encoding the stabilizers in tableaux that can be efficiently simulated.
@@
Figure 11: Time evolution for the $D_{4}$ model on three $30\times 30$ sized interpenetrating honeycomb lattices at three representative points in the $\gamma_{x}-\gamma_{z}$ phase diagram (see Fig. 1(b)). Plots show the densities of $X$ and $Z$ flags, stabilizer defect densities $n^{d}_{B}$ and $n^{d}_{A}$, and the logical error rate $1-F_{|\widetilde{0}\rangle_{L}(|\widetilde{+}\rangle_{L})}$ (see Eq. (42)) as a function of time. Time is measured in units of inverse erasure noise rate, set to $\eta=1$. The flags are initialized to $n^{f}_{X}=n^{f}_{Z}=0$ and the initial state of qubits is set to $|\widetilde{0}\rangle_{L}$, except when calculating $1-F_{|\widetilde{+}\rangle_{L}}$, where it is initialized in the $|\widetilde{+}\rangle_{L}$ state. Each data point is obtained by averaging over $500$ independent Monte Carlo trajectories. (a) Absorbing phase ($\gamma_{x}=4,\gamma_{z}=500$): when the $X$ correction rate is small, $X$ flags proliferate, which also renders $Z$ correction ineffective (see Eqs. (32), (33)). Since all flags are trapped in the absorbing state, the defect densities saturate to $1/2$, and all logical information is lost, as witnessed by $F\rightarrow 0$. (b) Partially $X$-active phase ($\gamma_{x}=14,\gamma_{z}=500$): while the densities of $X$ flags and vertex defects saturate to small values, $Z$ flags still proliferate. This phase can only protect classical information, as seen by perfect recovery of $|\widetilde{0}\rangle_{L}$ state (magenta curve) but failure to preserve $|\widetilde{+}\rangle_{L}$ state (cyan curve). (c) Active phase ($\gamma_{x}=35,\gamma_{z}=500$): when the $X$ correction rate is sufficiently large, the long-time value of $n^{f}_{X}$ (and of $n^{d}_{B}$) becomes small, allowing $Z$ correction to become sufficiently effective to stabilize $n^{f}_{Z}$ and $n^{d}_{A}$ densities at small values. In this phase, the logical error probability is close to 0, witnessed by $(1-F)\approx 0$ for both $|\widetilde{0}\rangle_{L}$ and $|\widetilde{+}\rangle_{L}$ initial states leading to protected quantum information up to long times.
@@ V.1 Dynamical phase diagram of flags
Fig. 11 shows the combined evolution of flag densities, stabilizer defect densities, and logical fidelity at a representative parameter value within each distinct steady-state phase shown in Fig. 1(b). We see that in each case, the dynamics of stabilizer defects closely tracks that of the flags, with both steady-state values and the time-scales at which these reach their long-time values closely correlated. Here, the densities of the vertex and plaquette stabilizer defects are defined, respectively, as
@@ V.2 Diagnostics for topologically ordered steady-states
Let us begin with the error-correction metric. The simultaneous +1+1 eigenspace of the $A_{p}$ and $B_{v}$ operators of the $D_{4}$ model is $22$-fold degenerate on the torus, reflecting the number of distinct superselection sectors (anyons) of $D_{4}$ TO. The states in this subspace can be distinguished based on the values of certain non-contractible loop operators. One choice is to label the $22$ ground states according to their eigenvalues under the non-contractible Abelian loop operators, given by $\mathcal{Z}^{c}_{H/V}=\prod_{k\in l_{H/V}}Z_{k}=\pm 1$ where $l_{H}(l_{V})$ are horizontal (vertical) loops on the dual of the $c$ colored honeycomb lattice, as shown in Fig 12. We term these the “logical-ZZ” operators. While there are $2^{6}$ different choices for the eigenvalues of these logical operators, only $22$ of them are consistent with $A_{p}=B_{v}=1$ for all $p,v$ [67]. The operators that switch between the allowed logical-$Z$ eigenstates are the non-contractible string operators associated with transporting non-Abelian anyons around the torus. We refer to these as the “logical-XX” operators, $\mathcal{X}^{c}_{H},\mathcal{X}^{c}_{V}$. As shown in Fig 12, these non-Abelian string operators consist of a string of $X$ operators decorated with non-local $CZ$ gates. A second basis choice for the $22$ locally indistinguishable ground states is by their values under the vertical $\mathcal{Z}_{V}$ and $\mathcal{X}_{V}$ logical operators.
In order to analyze the fate of the encoded logical information in the ground-space of the $D_{4}$ model under the noise and correction dynamics, it will be useful to consider two different initial logical states, one in each of these basis choices:
@@
Our second diagnostic utilizes the two-way channel connectivity equivalence relation placed on the space of mixed-states [55]: two states $\rho_{1}$ and $\rho_{2}$ belong to the same mixed-state phase if there exist a pair of quasi-local, finite-depth quantum channels $\Sigma_{12},\Sigma_{21}$ such that $\Sigma_{21}(\rho_{1})=\rho_{2}$ and $\Sigma_{12}(\rho_{2})=\rho_{1}$. In our case, the challenge is to determine whether there exists a finite-depth quantum channel that takes the steady-state density matrix $\rho_{ss}$ back to the initial error-free state $\rho_{1}$. We assess this by measuring the time required for the steady-state to evolve back to the error-free state under the Lindbladian Eq. (34) in the absence of noise i.e., under the action of only the error correction terms. Specifically, we examine how the time $t$ at which $\rho_{1}\sim e^{\mathcal{L}_{\eta=0}t}\rho_{ss}$ scales with system size, with $\mathcal{L}_{\eta=0}$ given by Eq. (34) with $\eta=0$. The density matrices $\rho_{1}$ and $\rho_{ss}$ are considered as being connected via a finite-depth quantum channel as long as this time grows at most (poly)-logarithmically with $L$ [55]. Here, two mixed states are considered to be close in terms of the fidelity, i.e. $\rho\sim\rho^{\prime}$, if $F(\rho,\rho^{\prime})>1-\varepsilon$ for some $\varepsilon>0$.
We emphasize that this is, in principle, a different (and stronger) criterion than requiring that the logical information encoded in the initial state can be recovered, since our MWPM decoder is highly non-local and need not guarantee the existence of a pair of quasi-local finite-depth two-way channels. Interestingly, however, at least in the context of the Toric code, Ref. [55] showed that the two criteria are equivalent in practice.